OpenStudy (anonymous):

A square has a perimeter of 12 cm. What is its area?

OpenStudy (zehanz):

 The square has four sides, each called s. Add them up. you get 12. Now calculate s. Then: how do you calculate the area of a square?

OpenStudy (anonymous):

i am still a little confused

OpenStudy (zehanz):

How many times s is 12?

OpenStudy (strawberry17):

Since the perimeter is 12 and each side is equal, you can just divide 12 by 4 to get each length side. For a square, to find the area, all you have to do is multiply s by s. The value of s is 4 (in this case).

OpenStudy (zehanz):

So s+s+s+s=12, or 4s=12, so s=12/4. We've found : s=3. So you have a square with side 3. Now how would you calculate the area?

OpenStudy (anonymous):

is the answer 36

OpenStudy (zehanz):

 Just count the little squares...

OpenStudy (anonymous):

9

OpenStudy (zehanz):

The area of a square with side s is defined as \(A=s^2\). Set s=3 in this formula. Or look at the second drawing :) Yes, it's 9!